Session: Feb 1998 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Sat Feb 28 15:57:05 MET 1998) Success
The experiment ran smoothly, without apparent problems. Tsys higher than it should be.

Onsala (A. Polatidis at Sat Feb 28 16:23:30 MET 1998) Success
No apparent problems, partially cloudy initially but fair weather after 10UT. Only HEAD 1 working hence no readback tests.

Jodrell (Alastair Gunn at Sun Mar 1 18:02:17 MET 1998) Minor failures
Missed start of experiment due to persistent high winds. Data good from about 05:11 UT. No other known problems.

Urumqi (Liu Xiang at Mon Mar 2 05:28:14 MET 1998) Minor failures
The front-end problem, so lost two scans the record began at UT:05:58, the Tsys is higher than normal value, we should measure it acurately later, clock offset is 2.52+/-0.05us.

Metsahovi (Kaj Wiik at Mon Mar 2 11:21:15 MET 1998) Success
No apparent problems. Some spurious error messages about head positioning before the start of the session.

Noto (C.Trigilio at Mon Mar 2 11:51:45 MET 1998) Success

Effelsberg (dgraham at Thu Mar 5 16:25:40 MET 1998) Minor failures
059-152 to 1344 elevation too low.